OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y x - 5 y -x - 4 (-5, 2) (5, -2) (-5, -2) (5, 2)

OpenStudy (anonymous):

\(\ge \) \(\le \) \(\ > \) \(\ < \) which inequality sign(s) are you using?

OpenStudy (anonymous):

=< and>=

OpenStudy (anonymous):

\[y \le x - 5\] \[y \ge -x - 4\]

OpenStudy (anonymous):

the first one is like this y=<x-5

OpenStudy (anonymous):

didnt i write it correct??

OpenStudy (anonymous):

ohh yea

OpenStudy (zzr0ck3r):

lets check (-5,2) y<=x-5 y>=−x−4 1st) 2<=-5-5 = -10 this already does not make sense, so no need to check the second equation. move on.. now try (5,-2)

OpenStudy (anonymous):

k still confused